how to fast splve of this equation:
(x+y)^3 this equation ans is this x^3+y^3+3xy(x+y)
but many time this situation will be accure...
gives one equation and find that equation main formulas.
x^3+y^3+3xy(x+y) this equation ans is you know that . (x+y)^3 but how.????

Answer by stanbon(75887)

x^3+y^3+3xy(x+y) this equation ans is you know that . (x+y)^3 but how.????
----
You have to learn to recognize the pattern.
(x+y)^3 = 3C3x^3 + 3C2x^2y + 3C1xy^2 + 3C0y^3
----
= x^3 + 3x^2y + 3xy^2 + y^3
===============================
Your x^3 + y^3 + 3xy(x+y) = x^3 + y^3 + 3x^2y + 3xy^2
Rearrange in decending powers of "x" and you get: x^3 + 3x^2y + 3xy^2 + y^3
which is (x+y)^3
==========================
Cheers,
Stan H